# Disentangled and Self-Explainable Node Representation Learning

### v1

@@ 3.2 Our approach: DiSeNE
Building upon the above components, introduced to satisfy our desiderata for interpretable node embeddings, we present our approach, DiSeNE. Specifically, DiSeNE takes as input the raw node attributes $\mathbf{X}\in\mathbb{R}^{|\mathcal{V}|\times F}$ and, depending on the encoder architecture, also the adjacency matrix $\mathbf{A}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$. The input is encoded into an intermediate embedding layer $\mathbf{Z}\in\mathbb{R}^{|\mathcal{V}|\times D}$. Next, DiSeNE processes the embedding matrix $\mathbf{Z}$ to compute the likelihood of link formation between node pairs, given by $\hat{y}(u,v;\mathbf{h})=\sigma(\mathbf{h}(u)^{\top}\mathbf{h}(v))$ where $\mathbf{h}(v)=\rho(\mathbf{W}^{\top}\mathbf{z}(v))$ are the final node representations in $\mathbf{H}\in\mathbb{R}^{|\mathcal{V}|\times K}$, obtained by applying a linear transformation $\mathbf{W}\in\mathbb{R}^{D\times K}$ followed by a non-linear activation function $\rho$. To encode $\mathbf{z}$, we employ architectures incorporating fully connected layers and graph convolutional layers [WSZ+19]. This process can be further enhanced by integrating more complex message-passing mechanisms or MLP operations. For example, the message-passing could initiate from an MLP-transformed node attribute matrix, MLP($\mathbf{X}$), or incorporate more sophisticated architectures beyond simple graph convolutions for increased expressiveness [XHLJ18, VCC+17].
The embeddings are optimized by combining the previously described objective functions for preserving structural faithfulness and achieving structural disentanglement, thereby improving dimensional interpretability. To avoid degenerate disentanglement solutions–specifically, the emergence of “empty” clusters characterized by near-zero columns in $\mathbf{F}$ that, while orthogonal to others, fail to convey meaningful information–we introduce a regularization strategy. This regularization ensures a minimal but significant level of connectivity within each topological substructure. Specifically, we enforce that the total amount of predicted edges in each anchor subgraph $\mathcal{G}_{k}$ $\sum_{u,v\in\mathcal{V}}\phi_{k}(u,v;\mathbf{h})$ to be non-zero. We found a more stable and precise approach by enforcing that the aggregated node features of each embedding dimension are non-zero, achieved by maximizing the entropy:
$\mathcal{H}=-\sum_{d=1}^{K}\left(\frac{\sum_{u}h_{d}(u)}{||\sum_{u}\mathbf{h}(u)||_{1}}\right)\log\left(\frac{\sum_{u}h_{d}(u)}{||\sum_{u}\mathbf{h}(u)||_{1}}\right).$
Thus, the model is optimized by minimizing the following comprehensive loss function:
@@
The hyperparameter $\lambda_{\text{ent}}$ determines the strength of the regularization, controlling the stability for explanation subgraph sizes across the various latent dimensions.
@@ 3.3 Proposed Evaluation Metrics
In the following, we introduce novel metrics to quantify multiple aspects related to interpretability and disentanglement in node embeddings, which we use to compare models in our experiments.
Comprehensibility.
@@
Comprehensibility measures how closely the identified topological explanations align with ground-truth clusters, which are crucial in the evolution of numerous complex real-world systems, i.e. community modules [GN02, HDF14]. We evaluate comprehensibility by treating edges in explanation masks $\{\mathbf{M}^{(d)}\}_{d=1,\dots,K}$ as retrieved items from a query, and measuring their overlap with the edges in the ground-truth communities using precision, recall, and $F_{1}$-score. Let $\mathcal{C}(\mathcal{E})=\{\mathcal{C}^{(1)},\dots,\mathcal{C}^{(m)}\}$ denotes the set of truthful link communities of the input graph222 Synthetic graphs can be constructed with ground-truth relevant sub-structures (like BA-Shapes [YBY+19] or SBM graphs). In real-world graphs, it is usually reasonable to assume that the community structure [For10] can serve as ground-truth.. Associated to partition $\mathcal{C}^{(i)}$, we define ground-truth edge masks $\mathbf{C}^{(i)}\in\{0,1\}^{V\times V}$ with binary entries $C^{(i)}_{uv}=[(u,v)\in\mathcal{C}_{i}]$. Comprehensibility score is given by the maximum $F_{1}$-score across ground-truth index:
$\mathrm{Comp}(\mathbf{M}^{(d)})=\mathrm{max}_{i}\left\{F_{1}(\mathbf{M}^{(d)},\mathbf{C}^{(i)})\right\}=\mathrm{max}_{i}\left\{\frac{2}{\mathrm{prec}(\mathbf{M}^{(d)},\mathbf{C}^{(i)})^{-1}+\mathrm{rec}(\mathbf{M}^{(d)},\mathbf{C}^{(i)})^{-1}}\right\}$ (5)
For precision, we weigh relevant item scores with normalized embedding masks values: $\mathrm{prec}(\mathbf{M}^{(d)},\mathbf{C}^{(i)})=\frac{\sum_{u,v}M^{(d)}_{uv}C^{(i)}_{uv}}{\sum_{u,v}M^{(d)}_{uv}}$. For recall, we weigh binarized embedding masks values with normalized ground-truth scores [^3] For the precision, we normalize with the sum of scores because they are continuous. For recall, we use the cardinality in place of the sum because the ground-truth has binary scores.: $\mathrm{rec}(\mathbf{M}^{(d)},\mathbf{C}^{(i)})=\frac{\sum_{u,v}[M^{(d)}_{uv}>0]C^{(i)}_{uv}}{|\mathcal{C}^{(i)}|}$.
@@
We refer to sparsity as the amount of disorder in the dimension’s explanations, it is generally defined as the ratio of the number of bits needed to encode an explanation compared to those required to encode the input [FKRA22]. Given that concise explanations are more effective in delivering clear insights, enhancing human understanding, we evaluate sparsity by measuring the normalized Shannon entropy over the mask distribution:
$\mathrm{Sp}(\mathbf{M}_{d})=-\frac{1}{\log|\mathcal{E}|}\sum_{(u,v)\in\mathcal{E}}\left(\frac{M^{(d)}_{uv}}{\sum_{u^{\prime},v^{\prime}}M^{(d)}_{u^{\prime}v^{\prime}}}\right)\log\left(\frac{M^{(d)}_{uv}}{\sum_{u^{\prime},v^{\prime}}M^{(d)}_{u^{\prime}v^{\prime}}}\right).$ (6)
A lower entropy in the mask distribution indicates higher sparsity.
@@
In explaining latent space representations, it is essential to comprehend how input space factors influence specific latent features. A well-structured, disentangled latent space should correspond to distinct, uncorrelated topological structures. We aim to quantify how different topological components affect pairwise feature correlations in the latent space. To achieve this, we propose a metric that measures the strength of association between the physical overlap of the explanation substructures $\{\mathcal{G}_{d}\}$ and the correlation among corresponding latent dimensions $\{\mathbf{H}_{:,d}\}$. We compute the overlap between two subgraph components using the Jaccard Similarity Index (JSI) of their edge sets from Eq. (2): $\mathrm{JSI}(d,l;\mathbf{h})=\frac{|\mathcal{E}_{d}\cap\mathcal{E}_{l}|}{|\mathcal{E}d\cup\mathcal{E}_{l}|}$. The overlap consistency (OvC) metric measures the linear correlation between the pairwise JSI values and the squared Pearson correlation coefficients ($\rho^{2}$) of the embedding features:
$\mathrm{OvC}(\mathbf{h})=\rho\Big{(}[\mathrm{JSI}(d,l;\mathbf{h})]_{d<l},[\rho^{2}(\mathbf{H}_{:,d},\mathbf{H}_{:,l})]_{d<l}\Big{)}$ (7)
where $[*]_{d<l}$ denotes the condensed list of pair-wise similarities. By using $\rho^{2}$ we remain agnostic about the sign of the correlation among latent features, since high overlaps could originate from both cases. This metric provides a systematic way to assess the extent to which topological partitions align with the distribution of features in the embedding space, thus offering deeper insights into the interpretability and disentanglement of the learned representations.
@@
In unsupervised node embeddings, it is crucial to assess how well the latent space captures the graph’s structure by encoding the positional relationships of nodes. An effective representation should preserve meaningful spatial properties that reflect node proximity and connectivity patterns. To achieve this, we propose to measure the extent to which node entries reflect their positions within the graph. Typically, positional encoding [LCSH21, LWWL20, YYL19] involves the use of several sets of node anchors $\mathcal{S}_{d}\subset\mathcal{V}$ that establish an intrinsic coordinate system. This system influences the node $u$’s features based on the node’s proximity $\zeta(u,\mathcal{S}_{d})=\mathrm{AGG}(\{\zeta(u,v),v\in\mathcal{S}_{d}\})$, where $\mathrm{AGG}$ denotes a specific pooling operation. As node proximity, we used the inverse of the shortest path distance $\zeta_{spd}(u,v)\equiv{(1+d_{spd}(u,v))}^{-1}$. As the anchor sets, we chose the embedding substructures used for explanations, $\mathcal{S}_{d}\equiv\mathcal{V}_{d}$.
For a specified pair of dimensions $(d,l)$, we assess the correlation between node features along dimension $d$ and the corresponding distances to the topological component indexed by $l$ via Feature-Proximity Correlation: $\mathrm{FPC}(d,l;\mathbf{h})=\rho\Big{(}\big{[}\zeta_{spd}(u,\mathcal{V}_{d})\big{]}_{u\in\mathcal{V}},\mathbf{H}_{:,l}\Big{)}$. The positional coherence metric (PoC) is defined to specifically evaluate the degree to which each feature $d$ is uniquely correlated with its corresponding topological component $\mathcal{V}_{d}$, without being significantly influenced by correlations with other substructures. This metric is calculated as the ratio of the average FPC for the given dimensions to the average FPC computed with pairs of permuted dimensions:
@@
$\mathrm{PoC}(\mathbf{h})=\frac{\sum_{d}\mathrm{FPC}(d,d;\mathbf{h})}{\bigl{\langle}\sum_{d}\mathrm{FPC}(d,\pi(d);\mathbf{h})\bigr{\rangle}_{\pi}}$ (8)
@@
where .$\langle.\rangle_{\pi}$ denotes an empirical average over multiple permutations. By comparing with random feature-subgraph pairs, the metric avoid promoting models with redundancies in the latent features, where high correlations with other topological components are possible.
Plausibility
@@
Figure 4: Schematic view of Plausibility metric computation. A high plausibility score indicates that the dimensions deemed more comprehensible also received higher importance scores from the post-hoc feature attribution technique.
Given the importance of node representations in downstream tasks, it is crucial to assess whether the features influencing predictions align with human expectations using synthetic benchmarks. To this end, we construct instance-wise explanations to determine if key features correspond to the topological structures behind the ground-truth labels. Typical post-hoc explainers [BGG+23] that produce feature importance scores often fail in this context because node embeddings are inherently uninterpretable, leading to useless explanations. Our approach overcomes this limitation by mapping feature importance back to the graph’s structural components, enabling a more meaningful evaluation of how well the embeddings capture the underlying factors driving node behavior.
Consider training a downstream binary classifier $b:\mathbb{R}^{K}\rightarrow[0,1]$, such as for node classification or link prediction. We detail the procedure for link prediction here (see also Figure 4), but we will also report the methodology for node classification in Appendix A.4. For an edge instance $\mathbf{h}(u,v)$ (which could be derived from node-pair operations such as $\mathbf{h}(u)\odot\mathbf{h}(v)$), we employ post-hoc methods to determine the feature relevance for the classifier prediction on the node pair instance $(u,v)$, $\{\Psi_{j}(u,v;b)\}_{j=1,\dots,K}$ that outputs important score for each of the embedding dimensions. Similarly to scores in Eq. (1), we define task-based masks $\mathbf{B}^{(j)}\in\mathbb{R}^{V\times V}_{\geq 0}$ which aggregate the individual logics of the classifier predictions: $B^{(j)}_{uv}(\Psi_{j};b)=\mathrm{max}\{0,\Psi_{j}(u,v;b)\}.$ To evaluate the consistency of a prediction, we consider the $F_{1}$-score in Eq. (5) related to the ground-truth structure of the edge under study, indexed by $g(u,v)$. Specifically, we define plausibility for an individual prediction as the average comprehensibility relative to the instance ground-truth, weighted by the computed feature importance:
$P\ell(u,v;b)=\frac{\sum_{d=1}^{K}f(\Psi_{j}(u,v;b))F_{1}(\mathbf{B}^{(j)},\mathbf{C}^{g(u,v)})}{\sum_{j=1}^{K}f(\Psi_{j}(u,v;b))}$ (9)

### v2

@@ 3.2 Our Approach: DiSeNE
Building upon the above components, introduced to satisfy our desiderata for interpretable node embeddings, we present our approach, DiSeNE. Specifically, DiSeNE takes as input identity matrix $\mathbb{\mathbf{1}}_{|\mathcal{V}|\times|\mathcal{V}|}$ as node attributes and, depending on the encoder architecture, also the adjacency matrix $\mathbf{A}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$. The input is encoded into an intermediate embedding layer $\mathbf{Z}\in\mathbb{R}^{|\mathcal{V}|\times D}$. Next, DiSeNE processes the embedding matrix $\mathbf{Z}$ to compute the likelihood of link formation between node pairs, given by $\hat{y}(u,v;\mathbf{h})=\sigma(\mathbf{h}(u)^{\top}\mathbf{h}(v))$ where $\mathbf{h}(v)=\rho(\mathbf{W}^{\top}\mathbf{z}(v))$ are the final node representations in $\mathbf{H}\in\mathbb{R}^{|\mathcal{V}|\times K}$, obtained by applying a linear transformation $\mathbf{W}\in\mathbb{R}^{D\times K}$ followed by a non-linear activation function $\rho$. To encode $\mathbf{z}$, we employ architectures incorporating fully connected layers and graph convolutional layers (Wu et al., 2019). This process can be further enhanced by integrating more complex message-passing mechanisms or MLP operations. For example, the message-passing could initiate from an MLP-transformed node attribute matrix, MLP($\mathbf{X}$), or incorporate more sophisticated architectures beyond simple graph convolutions for increased expressiveness (Xu et al., 2019; Velickovic et al., 2017).
[…]
@@
The hyperparameter $\lambda_{\text{ent}}$ determines the strength of the regularization, controlling the stability for explanation subgraph sizes across the various latent dimensions. We report the pseudo-code of DiSeNE in Appendix B, along with a space-time complexity analysis, showing that our method has $\mathcal{O}(|\mathcal{E}|K+|\mathcal{V}|K^{2}+|\mathcal{V}|KTL)$ runtime complexity and $\mathcal{O}(|\mathcal{V}|K+|\mathcal{V}|L)$ space complexity ($T$ refers to the window size, $L$ to simulated walks length), which is in line with well-established techniques for node embeddings (Tsitsulin et al., 2021). Our approach deviates from typical GNNs by focusing solely on learning from graph topology, as node attributes may not always align with structural information (e.g., in cases of heterophily (Zhu et al., 2024)). While semantic features can be integrated in various ways (Tan et al., 2024), we chose a more straightforward, broadly applicable method. Our node embeddings produce interpretable structural features $\mathbf{h}(u)$, which can be concatenated with node semantic attributes $\mathbf{x}(u)$, enabling transparent feature sets for effective explanations in downstream tasks via post-hoc tabular techniques.
@@ 3.3 Proposed Evaluation Metrics
In the following, we present novel metrics to quantify interpretability and disentanglement in unsupervised node embeddings, which we use to compare models in our experiments. Unlike prior works, which primarily focus on explaining graph model decisions, our approach offers a novel perspective by targeting the explanation of graph model encodings. Traditional graph-based explainers focus on interpreting GNN model decisions for specific tasks, highlighting subgraphs or motifs responsible for the predictions (Longa et al., 2025; Li et al., 2025). In contrast, our objective is to assess the interpretability of GNN models in a task-agnostic setting, by evaluating explanatory subgraphs derived from node embedding dimensions that are learned without supervision. In the next, we define our comprehensive vocabulary of evaluation metrics. While alternative taxonomies exist in the literature, we propose a consistent framework highlighting similarities and differences between our indicators and other established metrics in the field of graph-based explainability (Yuan et al., 2023; Kakkad et al., 2023).
Topological Alignment.
This metric measures how well the explanations of embedding dimensions align with human-interpretable graph structures. Given their importance in the organization of complex real-world systems (Girvan & Newman, 2002; Hric et al., 2014), community modules (or clusters) are often the most intuitive units for understanding a graph. We evaluate topological alignment by handling edges in explanation masks $\{\mathbf{M}^{(d)}\}_{d=1,\dots,K}$ as retrieved items from a query, and measuring their overlap with the edges in the ground-truth communities using precision, recall, and $F_{1}$-score. While this metric captures one aspect of human-comprehensibility, it does not exhaust the space of interpretable structures, such as motifs, roles, or domain-specific patterns. For example, in biological networks, structures like protein complexes or regulatory circuits are often considered as more interpretable. In social networks, roles like hubs, bridges, or peripheral nodes may give better explanations than just communities. Future work could incorporate alternative structural annotations to evaluate broader forms of topological alignment.
@@
Let $\mathcal{C}(\mathcal{E})=\{\mathcal{C}^{(1)},\dots,\mathcal{C}^{(m)}\}$ denotes the set of truthful link communities of the input graph [^3] Synthetic graphs can be constructed with ground-truth relevant sub-structures (like BA-Shapes (Ying et al., 2019) or SBM graphs). In real-world graphs, it is usually reasonable to assume that the community structure (Fortunato, 2010) can serve as ground-truth.. Associated to partition $\mathcal{C}^{(i)}$, we define ground-truth edge masks $\mathbf{C}^{(i)}\in\{0,1\}^{V\times V}$ with binary entries $C^{(i)}_{uv}=\mathbb{1}[(u,v)\in\mathcal{C}^{(i)}]$. For a given mask $\mathbf{M}^{(d)}$, topological alignment score is given by the maximum edge overlap (as $F_{1}$-score) of the explanation substructure computed across community structures in $\mathcal{C}(\mathcal{E})$, quantifying how closely the explanation match human-understandable ground-truth:
$\mathrm{Align}(\mathbf{M}^{(d)})=\mathrm{max}_{i}\left\{F_{1}(\mathbf{M}^{(d)},\mathbf{C}^{(i)})\right\}=\mathrm{max}_{i}\left\{\frac{2}{\mathrm{prec}(\mathbf{M}^{(d)},\mathbf{C}^{(i)})^{-1}+\mathrm{rec}(\mathbf{M}^{(d)},\mathbf{C}^{(i)})^{-1}}\right\}$ (4)
For precision, we weigh relevant item scores with normalized embedding masks values: $\mathrm{prec}(\mathbf{M}^{(d)},\mathbf{C}^{(i)})=\frac{\sum_{u,v}M^{(d)}_{uv}C^{(i)}_{uv}}{\sum_{u,v}M^{(d)}_{uv}}$. For recall, we weigh binarized embedding masks values with normalized ground-truth scores [^4] For the precision, we normalize with the sum of scores because they are continuous. For recall, we use the cardinality in place of the sum because the ground-truth has binary scores.: $\mathrm{rec}(\mathbf{M}^{(d)},\mathbf{C}^{(i)})=\frac{\sum_{u,v}\mathbb{1}[M^{(d)}_{uv}>0]C^{(i)}_{uv}}{|\mathcal{C}^{(i)}|}$. This approach is similar to the accuracy assessment used in GNNExplainer and subsequent works (Ying et al., 2019; Luo et al., 2020), where explanations for model decisions are compared to planted graph substructures, perceived as human-readable justifications for a node’s prediction. However, instead of evaluating individual decisions’ explanations by using synthetic ground-truth motifs, we perform a global assessment of the unsupervised embedding dimensions by comparing their explanation subgraphs with a reasonable ground-truth about the graph structure. An analogous evaluation of explanation correctness which specifically focuses on individual model decisions will be provided later by the plausibility metric.
@@
We refer to sparsity as a measure of the localization of subgraph explanations, it is generally defined as the ratio of the number of bits needed to encode an explanation compared to those required to encode the input (Pope et al., 2019; Funke et al., 2023). As we produce soft masks, we use entropy for this quantification. Given that compact explanations are more effective in delivering clear insights, we evaluate sparsity by measuring the normalized Shannon entropy over the mask distribution:
$\mathrm{Sp}(\mathbf{M}^{(d)})=-\frac{1}{\log|\mathcal{E}|}\sum_{(u,v)\in\mathcal{E}}\left(\frac{M^{(d)}_{uv}}{\sum_{u^{\prime},v^{\prime}}M^{(d)}_{u^{\prime}v^{\prime}}}\right)\log\left(\frac{M^{(d)}_{uv}}{\sum_{u^{\prime},v^{\prime}}M^{(d)}_{u^{\prime}v^{\prime}}}\right)$. (5)
A lower entropy in the mask distribution indicates higher sparsity/compactness. Motivation for using entropy for explanation size quantification can be found in one of the existing works (Funke et al., 2023).
@@
This metric assesses whether the correlation between two embedding dimensions is mirrored in their corresponding explanations. A well-structured, disentangled latent space should correspond to distinct, uncorrelated topological structures. For example, if two embedding dimensions are correlated, their explanation substructures should also overlap. In our context, this a proxy measure for explanation’s contrastivity (Wang et al., 2023; Pope et al., 2019), based on the intuition that explanations for unrelated dimensions should differ substantially—particularly in the specific substructures they highlight. We aim to quantify how different topological components affect pairwise feature correlations in the latent space. To achieve this, we propose a metric that measures the strength of association between the physical overlap of the explanation substructures $\{\mathcal{G}_{d}\}$ and the correlation among corresponding latent dimensions $\{\mathbf{H}_{:,d}\}$. We compute the overlap between two subgraph components using the Jaccard similarity index of their edge sets from Eq. (2): $J(d,l;\mathbf{h})=\frac{|\mathcal{E}_{d}\cap\mathcal{E}_{l}|}{|\mathcal{E}d\cup\mathcal{E}_{l}|}$. The overlap consistency (OvC) metric measures the linear correlation between pairwise Jaccard values and squared Pearson correlation coefficients ($\rho^{2}$) of the embedding features:
$\mathrm{OvC}(\mathbf{h})=\rho\Big([J(d,l;\mathbf{h})]_{d<l},[\rho^{2}(\mathbf{H}_{:,d},\mathbf{H}_{:,l})]_{d<l}\Big)$ (6)
where $_{d<l}$ denotes the condensed list of pair-wise similarities. By using $\rho^{2}$ we remain agnostic about the sign of the correlation among latent features, since high overlaps could originate from both cases.
@@
This metric evaluates whether the feature value of a node representation in a specific embedding dimension corresponds to its spatial relationship with the explanation substructure for that dimension. In fact, an effective representation should preserve meaningful spatial properties that reflect node proximity and connectivity patterns. In our context, this is a proxy measure for explanation’s faithfulness (Zhao et al., 2023b), reflecting how well the dimension-based structures align with the embeddings they are intended to explain. To achieve this, we propose to measure the extent to which node entries reflect their relative positions with the subgraphs used as explanations. For example, if node $u$ has a high value in embedding dimension $d$, it should be positioned closer to the substructure that explains dimension $d$. Typically, positional encoding (Rampásek et al., 2022; Li et al., 2020; You et al., 2019) involves the use of several sets of node anchors $\mathcal{S}_{d}\subset\mathcal{V}$ that establish an intrinsic coordinate system. This system influences the node $u$’s features based on the node’s proximity $\zeta(u,\mathcal{S}_{d})=\mathrm{AGG}(\{\zeta(u,v),v\in\mathcal{S}_{d}\})$, where $\mathrm{AGG}$ denotes a specific pooling operation. As node proximity, we used the inverse of the shortest path distance $\zeta_{spd}(u,v)\equiv{(1+d_{spd}(u,v))}^{-1}$. As the anchor sets, we chose the embedding substructures used for explanations, $\mathcal{S}_{d}\equiv\mathcal{V}_{d}$. For a specified pair of dimensions $(d,l)$, we assess the correlation between node features along dimension $d$ and the corresponding distances to the topological component indexed by $l$ via feature-proximity correlation: $fp_{corr}(d,l;\mathbf{h})=\rho\Big(\big[\zeta_{spd}(u,\mathcal{V}_{d})\big]_{u\in\mathcal{V}},\mathbf{H}_{:,l}\Big)$. The positional coherence metric (PoC) is defined to specifically evaluate the degree to which each feature $d$ is uniquely correlated with its corresponding topological component $\mathcal{V}_{d}$, without being significantly influenced by correlations with other substructures. This metric is calculated as the ratio of the average $fp_{corr}$ for the given dimensions to the average $fp_{corr}$ computed with pairs of permuted dimensions:
@@
$\mathrm{PoC}(\mathbf{h})=\frac{\sum_{d}fp_{corr}(d,d;\mathbf{h})}{\bigl\langle\sum_{d}fp_{corr}(d,\pi(d);\mathbf{h})\bigr\rangle_{\pi}}$ (7)
@@
where $\langle.\rangle_{\pi}$ denotes an empirical average over multiple permutations. By comparing with random feature-subgraph pairs, the metric avoid promoting models with redundancies in the latent features, where high correlations with other topological components are possible.
Plausibility.
Plausibility evaluates how closely subgraph explanations align with human reasoning, a concept also studied in existing works like Bagel (Rathee et al., 2022). Unlike topological alignment, which measures the alignment of explanation substructures with community clusters in the graph (independent of task-specific information), plausibility focuses on explanations for decisions made in downstream tasks, using the embedding features as predictor variables. To this end, relying on feature-based post-hoc explanation methods (Lundberg & Lee, 2017; Ribeiro et al., 2016), we construct instance-specific explanations to determine feature importance scores related to the topological structures. Typical feature importance explainers (Bodria et al., 2023) are useless in this context because node embeddings have inherently uninterpretable features, leading to uninformative explanations. Our approach overcomes this limitation by mapping explanations back to the graph’s structural components, that are human-readable.
@@
Figure 4: Sketch of Plausibility metric computation. High plausibility scores indicate that the dimensions deemed more comprehensible also received higher importance scores from the post-hoc attribution technique.
We detail the procedure for link prediction here (see also Figure 4), but we also report the methodology for node classification in Appendix E. The process involves three key steps.
(1) Training a downstream classifier. We train a binary classifier $b:\mathbb{R}^{K}\rightarrow[0,1]$ to perform the downstream link prediction task using node embeddings. This classification model can either leverage feature interpretability techniques, such as SHAP (Lundberg & Lee, 2017), or being inherently interpretable, such as logistic regression.
(2) Extracting post-hoc explanations. For an edge instance $\mathbf{h}(u,v)$ (which could be derived from node-pair operations such as $\mathbf{h}(u)\odot\mathbf{h}(v)$), we employ post-hoc methods to determine the feature relevance for the classifier prediction on the node pair instance $(u,v)$, $\{\Psi_{j}(u,v;b)\}_{j=1,\dots,K}$ for each of the embedding dimensions. Similarly to Eq. (1) which defines task-agnostic masks, here we calculate task-specific masks $\mathbf{B}^{(j)}\in\mathbb{R}^{V\times V}_{\geq 0}$ which aggregate the logic of the classifier based on individual feature importance: $B^{(j)}_{uv}(\Psi_{j};b)=\mathrm{max}\{0,\Psi_{j}(u,v;b)\}.$ Substructures corresponding to the most important features (identified by the post-hoc techniques), or the explanations directly returned by the interpretable model serve as the final explanations.
(3) Evaluation. Explanations are then compared with the available human rationale behind the decisions. To do that, we resort the $F_{1}$-score in Eq. (4) with respect to the ground-truth structure of the edge under study, indexed by $g(u,v)$, that it is known has driven the classifier decision. Specifically, we define plausibility for an individual prediction $b(u,v)$ as the average $F_{1}$-accuracy relative to the ground-truth structure for the instance, weighted by the computed feature importance:
$P\ell(u,v;b)=\frac{\sum_{j=1}^{K}f(\Psi_{j}(u,v;b))F_{1}(\mathbf{B}^{(j)},\mathbf{C}^{g(u,v)})}{\sum_{j=1}^{K}f(\Psi_{j}(u,v;b))}$ (8)
@@
It is worth to note that although our measure is similar to the accuracy index used to assess correctness of explanation subgraphs (Ying et al., 2019), we refer to it as "plausibility." This is because, despite their competitive classification performance, typical embedding representations may fail to capture the truthful substructures in the most task-predictive features, as they are not constrained to follow human reasoning. Thus, with plausibility, we aim to understand when predictive embedding features are also congruent to human rationales. The choice of the weighting function $f$ is case-specific and tailored to the objective of the analysis. For example, choosing $f$ as indicator function such that $f(\Psi_{j})=\mathbb{1}[j=\mathrm{argmax}\{\Psi_{1},\dots\Psi_{K}\}]$ allows to study the plausibility of the most predictive substructure.